OpenStudy (anonymous):

Pathogirion therom If e = 9 cm and f = 40 cm, what is the length of g? 43 cm 31 cm 39 cm 41 cm

OpenStudy (whpalmer4):

That's Pythagorean theorem :-) If e and f are the shorter sides of the triangle, and this is a right triangle, then the length of the hypotenuse (g, presumably) will be related by \[e^2+f^2 =g^2\]If you solve that for \(g\), by taking the square root of both sides, \[g = \sqrt{e^2+f^2}\]Plug in your numbers and find the answer.